[Briana, Ian, Torrey]
We currently have the current controller limit set to 230 mA, but the power sent to the laser doesn't seem to be able to reach the standard 39.4 mW in the specs before the current applied reaches 250 mA. At 236 mA, temperature set at 9.99 kOhms (supposedly 25 deg Celsius), the power reading is 35 mW (with the 500 mW setting in place). This reading agrees with graphs on the spec sheet but when we set the current control limit, we should theoretically be able to turn up the current to some value <250 mA such that the maximum power output (55mW) is reached (see thorlabs video linked two posts back). The lower laser power may be because we aren't driving temperature high enough for the laser to reach the higher power output.
We did use the 5 mW setting on the photodetector at some point and measured 39.4 mW at 230 mA at temperature ~25 deg Celsius. We will not use the 5 mW setting moving forward because our power is much higher than that. The current/corresponding power readings on the previous post should be disregarded becaues we also used the 5 mW setting to measure things, which did not have the ND filter from the 500 mW setting in place.
We also set up the plan to shoot the 780 nm laser through the vapor cell and scan across the frequency to detect the Lamb dip (beam path in red in one of the pictures). This will be done with a photodiode (PDA36A2) connected to a Moku, which we will set up tomorrow. The temperature control for the vapor (TC300) has also been placed but not set up. We should also check the specs of the vapor cell to see what current is needed.
I misread the power sensor ND filter and am responsible for that mistake. I did not put enough power/area to get anywhere close to damaging the sensor. We should print a label for the sensor so in future people don't repeat my mistake.